Determine the Weight f the Sample, if it Used 24.9ml of 0.1Normal Sodium Hydroxide of Titrant {33.29% of Busulfan (C7H6O2)} 


1
Expert's answer
2022-03-15T06:56:03-0400

C7H6O2 is not the Busulfan. This is the benzoic acid.

M(C7H6O2) = 122 g/mole;

NaOH + C7H6O2 = C7H5O2Na + H2O;

n(NaOH) = C(NaOH) * V(NaOH) = 0.1 * 0.0249 = 0.00249 moles;

n(C7H6O2) = n(NaOH) = 0.00249 moles;

m(C7H6O2) = n(C7H6O2) * M(C7H6O2) = 0.00249 * 122 = 0,3 g;

m(sample) = m(C7H6O2) * 100 %/X(C7H6O2) = 0.3 * 100%/33.29% = 0.9 g.

Answer: 0.9 g.